OpenStudy (anonymous):

4/5divided by (-5/6)

myininaya (myininaya):

\[\frac{a}{b} \div \frac{c}{d}=\frac{a}{b} \cdot \frac{d}{c}\]

OpenStudy (anonymous):

-0.026666667

OpenStudy (anonymous):

its not 2/3?

myininaya (myininaya):

\[\frac{4}{5} \div \frac{-5}{6}=\frac{4}{5} \cdot \frac{-6}{5}\]

myininaya (myininaya):

what is the product of the top?

myininaya (myininaya):

what is the product of the bottom?

OpenStudy (anonymous):

4/5 times -6/5

myininaya (myininaya):

answer those two questions and you will have your answer

OpenStudy (anonymous):

do i just multiply across?

myininaya (myininaya):

find the product of the top then find the product of the bottom

OpenStudy (anonymous):

24/25?

myininaya (myininaya):

don't forget the negative sign and then yes that would be correct

OpenStudy (anonymous):

oh!!! okay!